OpenStudy (anonymous):

HELP!! solve sin^-1(√2/2) + cos^-1x = 2pi/3....do you start with solving by cos^-1x ?

OpenStudy (anonymous):

si the first step correct>

OpenStudy (anonymous):

i would start by knowing that \(\sin^{-1}(\frac{\sqrt{2}}{2})=\frac{\pi}{4}\)

OpenStudy (anonymous):

so start with \[\frac{\pi}{4}+\cos^{-1}(x)=\frac{2\pi}{3}\] then subtract \(\frac{\pi}{4}\) from both sides

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

wait why is it pi/4?

OpenStudy (anonymous):

oh sin45 is pi/4 okay

OpenStudy (anonymous):

\(\sin^{-1}(\frac{\sqrt{2}}{2})=\frac{\pi}{4}\) because \(\sin(\frac{\pi}{4})=\frac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

ignore degrees. you are clearly working in radians (numbers) because you have \(\frac{2\pi}{3}\)on the right

OpenStudy (anonymous):

the sine of 45 degrees is not \(\frac{\pi}{4}\)! if you measure an angle in degrees, then you could say \(\sin(45)=\frac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

okay. so i got up to here!

OpenStudy (anonymous):

pi/4 lies on 45 degrees i mean....is that correct?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

hello???

OpenStudy (anonymous):

too much work

OpenStudy (anonymous):

\[\frac{2\pi}{3}-\frac{\pi}{4}=\frac{5\pi}{12}\]

OpenStudy (anonymous):

oh i see, no not too much work !

OpenStudy (anonymous):

yes, you have it, i was thinking wrong

OpenStudy (anonymous):

last job is to plug in the numbers in what you wrote above and you will be done

OpenStudy (anonymous):

okay thank you!!

OpenStudy (anonymous):

yw